Cougars speed is 40 mph, and a zebras speed is 15 mph. How much time it takes the cougar to reach the zebra if the distance between the animals is 2.5 miles?
1
Expert's answer
2018-06-11T15:27:08-0400
The relative speed of the cougar with respect to zebra
V=v-u=40 mph-15 mph=25 mph
So, time needed for the cougar to reach a zebra
t=L/V=(2.5 mi)/(25 mph)=0.1 hr=6 min